### v2

@@ 4.1.3 Numerical (In)Stability of CR/MINRES
For the experiments in this section, we set $d=1,000$ and $m=800$. The maximum number of iterations is capped at $800$, and the termination condition is triggered at the threshold $10^{-8}$. To assess the extent of numerical inaccuracies, we monitor the fundamental theoretical properties of these methods, specifically orthogonality and conjugacy of the relevant vectors, as per Section 3-(b) and Section 3-(c). For CG, the deviation from the properties outlined in Section 3-(b) is calculated at each iteration. This involves gathering the update directions $\mathbf{p}_{k-1}$ and the normalized residual vectors $\mathbf{r}_{k-1}/\left\|\mathbf{r}_{k-1}\right\|$ as columns of matrices $\mathbf{P}_{k}$ and $\mathbf{R}_{k}$, respectively. The spectral norm of the matrices $\mathbf{I}_{k}-\mathbf{R}_{k}^{{\top}}\mathbf{R}_{k}$ and $\text{diag}(\{\left\langle\mathbf{p}_{i},\mathbf{Ap}_{i}\right\rangle\}_{i=0}^{i=k})-\mathbf{P}_{k}^{{\top}}\mathbf{AP}_{k}$ is then calculated, where “diag” refers to a diagonal matrix constructed from the given set of values. Following the properties outlined in Section 3-(c), similar quantities are also calculated for CR and MINRES, namely $\mathbf{I}_{k}-(\mathbf{AP}_{k})^{{\top}}\mathbf{AP}_{k}$ (here the vectors $\mathbf{Ap}_{k-1}$ are normalized) and $\text{diag}(\{\left\langle\mathbf{r}_{i},\mathbf{Ar}_{i}\right\rangle\}_{i=0}^{i=k})-\mathbf{R}_{k}^{{\top}}\mathbf{AR}_{k}$. In the absence of numerical issues, these quantities should all be evaluated to zero.
@@
In addition, we also monitor $\|(\mathbf{b}-\mathbf{Ax}_{k})-\mathbf{r}_{k}\|$ where $\mathbf{r}_{k}$ is the residual vector calculated naturally as part of the algorithm, for example, Line 5 in Algorithms 1 and 2 (a similar quantity also arise as part of MINRES; see 19). Theoretically, the residual calculated within the algorithm must align with that calculated from the iterate as $\mathbf{b}-\mathbf{Ax}_{k}$. In practice, however, since they are obtained through different means, we expect some discrepancies between them due to numerical errors. Nevertheless, methods that are inherently more stable should exhibit less severe discrepancies.
@@
The results of the experiments for this section are gathered Figures 7, 8 and 9. In stark contrast to CR and MINRES, CG exhibits a significant degree of numerical instability in Figure 7, reaching a level that can be categorized as catastrophic. These observations align well with the findings in 24 who suggests that CG under inconsistent systems could lead to divergence, though it does not provide any more insights. Furthermore, in Figures 8 and 9, although all methods experience some loss of orthogonality and conjugacy in their underlying vectors, CR and MINRES demonstrate slightly more stable numerical behavior. This is evident in the quantity $\|(\mathbf{b}-\mathbf{Ax}_{k})-\mathbf{r}_{k}\|$, and they make relatively satisfactory progress toward obtaining a normal solution with a relatively small residual from the normal equations.
@@ 4.2.1 Poisson Equation with Neumann Boundary Condition
where $u:\Omega\subseteq\mathbb{R}^{2}\to\mathbb{R}$ is the unknown function. Here, $\Delta u$, $\nabla u$ and $\mathbf{n}$ are the Laplace operator of $u$, the gradient of $u$, and the normal vector to the boundary $\partial\Omega$ (the exterior surface), respectively. We use a five-point stencil on uniform staggered grids to formulate the finite difference of the Laplace operator of $u$. The resulting Laplacian matrix of this particular problem is known to be PSD and rank deficient 5, 35.
@@ 4.2.2 Gaussian Smoothing
Next, we consider the behavior of CG, CR, and MINRES in the context of an image deblurring problem 36. For a given image $\mathbf{x}$, we apply a linear Gaussian smoothing / blurring operator $\mathbf{A}$, as a 2D convolutional operator, on the image and add some noise $\mathbf{e}$ sampled from a uniform distribution, i.e., $\mathbf{b}=\mathbf{A}\mathbf{x}+\mathbf{e}$. We follow a similar experimental setup as 34. The true image and blurry version are given in Figure 13. We remind that though $\mathbf{A}$ typically has full rank, it often contains several small singular values, which make it almost numerically singular. We do not use any particular regularization and simply apply CG, CR and MINRES to attempt to recover the original image. The maximum number of iterations and the termination condition are set to $300$ and $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\leq 10^{-5}$, respectively.
The results are presented in Figures 14 and 15. Unlike CR and MINRES, as evident from Figures 14 and 15, CG fails to achieve the required termination threshold before reaching the maximum number of iterations. Clearly, the image recovered by CG at termination is of significantly lower quality compared to those obtained by CR and MINRES. Similar to the previous section, Figure 15 showcases the best performance of CG — the image recovered by CG at the iteration with the smallest relative residual — compared with the images from CR and MINRES at the same iteration. At this particular iteration, while CG produces a slightly improved image, it remains visibly blurry with a larger relative error than those obtained by CR/MINRES at termination.
@@ 4.2.3 Kernel Regression
Finally, we consider kernel regression 37, which appears frequently in machine learning applications. Kernel regression is an example of non-parametric regression where given training data $\mathcal{D}=\{\mathbf{a}_{i},b_{i}\}_{i=1}^{n}\subset\mathbb{R}^{p}\times\mathbb{R}$, we aim to find a prediction function $f:\mathbb{R}^{p}\to$ that is a solution to the following minimization problem
@@
where $\mathcal{H}$ is some appropriate reproducing kernel Hilbert space (RKHS) with the kernel function $\kappa:^{p}\times^{p}\to$ 38. It can be shown that any solution to the above problem can be written as $f=g+h$ where $g=\sum_{j=1}^{n}x_{j}\kappa(\mathbf{a}_{j},.)$ and $h\perp\text{Span}\{\kappa(\mathbf{a}_{j},.)\}_{j=1}^{n}$, leading to the least-squares regression problem
@@
We use Wine Quality dataset, comprising $6,497$ training examples with $p=10$ 39. The dataset is randomly split into $80\%$ training (thus, $n=5,197$) and $20\%$ validation sets. Employing the radial basis function kernel $\kappa(\mathbf{a},\mathbf{a}^{\prime})=\exp{(-L\|\mathbf{a}-\mathbf{a}^{\prime}\|_{2}^{2})}$ with $L=10^{-4}$ results in the positive semi-definite kernel matrix $\mathbf{A}$ of rank $438$. The maximum number of iterations and the termination condition are set to $2,000$ and $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\leq 10^{-7}$, respectively. To assess the predictive performance of the recovered model, we measure the mean squared error (MSE) for the test set at each iteration. The results, depicted in Figure 16, clearly show that both CR and MINRES greatly outperform CG, not only in metrics related to the training set but also in terms of predictive performance as measured by the MSE on the test set.
@@ Appendix A Unifying Framework
In this section, we present some technical results for Algorithms 1 and 2 that serve as foundational tools in all of our derivations in this paper. To simplify the exposition, we adopt a more general framework depicted in Algorithm 5, which encapsulates both Algorithms 1 and 2 at the same time. Specifically, considering $s=0$ and $s=1$ in Algorithm 5 amount to Algorithm 1 and Algorithm 2, respectively. In essence, it is shown that all common properties of CG and CR, well-known for positive definite settings, continue to hold under indefinite matrices. The proofs are straightforward adaptation of the PD, found in many classical resources, e.g., 2, 40.
@@
for some $i$. We show the claim for $i+1$. Since $\mathbf{p}_{i-1}\in\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i-1}\mathbf{r}_{0})$, we have $\mathbf{Ap}_{i-1}\in\text{Span}(\mathbf{Ar}_{0},\mathbf{A}^{2}\mathbf{r}_{0},\cdots,\mathbf{A}^{i}\mathbf{r}_{0})$, which from $\mathbf{r}_{i}=\mathbf{r}_{i-1}-\alpha_{i-1}^{(s)}\mathbf{Ap}_{i-1}$ implies $\mathbf{r}_{i}\in\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i}\mathbf{r}_{0})$. So, $\text{Span}(\mathbf{r}_{0},\mathbf{r}_{1},\cdots,\mathbf{r}_{i})\subset\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i}\mathbf{r}_{0})$. However, from $\mathbf{A}^{i}\mathbf{r}_{0}=\mathbf{A}(\mathbf{A}^{i-1}\mathbf{r}_{0})$, we have
@@
by 9a and 10b and our induction assumptions.
Lemma A.5.
@@
where the first term and the second term each evaluate to zero by the inductive hypothesis and 9b, respectively. This gives us 11a. Now, let $j=k$, for $i=k$, we have $\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{p}_{k}\right\rangle=\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\left(\mathbf{r}_{k}+\beta_{k-1}^{(s)}\mathbf{p}_{k-1}\right)\right\rangle=\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle$, by 11a. For $i<k$, we have
@@
where the last line is by the definition of $\beta_{0}^{(s)}$, and the second equality due to 9a and the inductive hypothesis, respectively. This gives us 11b.
@@ Appendix B Further Results
Lemma B.1.
Under Algorithm 2, in Algorithm 5 we have $(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{p}_{k-1}=c^{(s)}_{k-1}\mathbf{b}^{\perp}$ for $1\leq k\leq g$, where $c^{(s)}_{k-1}=1+\beta^{(s)}_{k-2}c^{(s)}_{k-2},\;k\geq 1$, with $c^{(s)}_{0}=1$.
Proof B.2.

### v4

@@ 4.1.3 Numerical (In)Stability of CR/MINRES
For the experiments in this section, we set $d=1,000$ and $m=800$. The maximum number of iterations is capped at $800$, and the termination condition is triggered at the threshold $10^{-8}$. To assess the extent of numerical inaccuracies, we monitor the fundamental theoretical properties of these methods, specifically orthogonality and conjugacy of the relevant vectors, as per Property 1-(b) and Property 1-(c). For CG, the deviation from the properties outlined in Property 1-(b) is calculated at each iteration. This involves gathering the update directions $\mathbf{p}_{k-1}$ and the normalized residual vectors $\mathbf{r}_{k-1}/\left\|\mathbf{r}_{k-1}\right\|$ as columns of matrices $\mathbf{P}_{k}$ and $\mathbf{R}_{k}$, respectively. The spectral norm of the matrices $\mathbf{I}_{k}-\mathbf{R}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{R}_{k}$ and $\text{diag}(\{\left\langle\mathbf{p}_{i},\mathbf{Ap}_{i}\right\rangle\}_{i=0}^{i=k})-\mathbf{P}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{AP}_{k}$ is then calculated, where “diag” refers to a diagonal matrix constructed from the given set of values. Following the properties outlined in Property 1-(c), similar quantities are also calculated for CR and MINRES, namely $\mathbf{I}_{k}-(\mathbf{AP}_{k})^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{AP}_{k}$ (here the vectors $\mathbf{Ap}_{k-1}$ are normalized) and $\text{diag}(\{\left\langle\mathbf{r}_{i},\mathbf{Ar}_{i}\right\rangle\}_{i=0}^{i=k})-\mathbf{R}_{k}^{{{\mathchoice{\raisebox{0.0pt}{$\displaystyle\intercal$}}{\raisebox{0.0pt}{$\textstyle\intercal$}}{\raisebox{0.0pt}{$\scriptstyle\intercal$}}{\raisebox{0.0pt}{$\scriptscriptstyle\intercal$}}}}}\mathbf{AR}_{k}$. In the absence of numerical issues, these quantities should all be evaluated to zero.
@@
In addition, we also monitor $\|(\mathbf{b}-\mathbf{Ax}_{k})-\mathbf{r}_{k}\|$ where $\mathbf{r}_{k}$ is the residual vector calculated naturally as part of the algorithm, for example, Line 5 in Algorithms 1 and 2 (a similar quantity also arise as part of MINRES; see [28]). Theoretically, the residual calculated within the algorithm must align with that calculated from the iterate as $\mathbf{b}-\mathbf{Ax}_{k}$. In practice, however, since they are obtained through different means, we expect some discrepancies between them due to numerical errors. Nevertheless, methods that are inherently more stable should exhibit less severe discrepancies.
@@
The results of the experiments for this section are gathered Figures 7, 8 and 9. In stark contrast to CR and MINRES, CG exhibits a significant degree of numerical instability in Figure 7, reaching a level that can be categorized as catastrophic. These observations align well with the findings in [23] who suggests that CG under inconsistent systems could lead to divergence, though it does not provide any more insights. Furthermore, in Figures 8 and 9, although all methods experience some loss of orthogonality and conjugacy in their underlying vectors, CR and MINRES demonstrate slightly more stable numerical behavior. This is evident in the quantity $\|(\mathbf{b}-\mathbf{Ax}_{k})-\mathbf{r}_{k}\|$, and they make relatively satisfactory progress toward obtaining a normal solution with a relatively small residual from the normal equations.
@@ 4.2.1 Poisson Equation with Neumann Boundary Condition
where $u:\Omega\subseteq\mathbb{R}^{2}\to\mathbb{R}$ is the unknown function. Here, $\Delta u$, $\nabla u$ and $\mathbf{n}$ are the Laplace operator of $u$, the gradient of $u$, and the normal vector to the boundary $\partial\Omega$ (the exterior surface), respectively. We use a five-point stencil on uniform staggered grids to formulate the finite difference of the Laplace operator of $u$. The resulting Laplacian matrix of this particular problem is known to be PSD and rank deficient [44, 37].
@@ 4.2.2 Gaussian Smoothing
Next, we consider the behavior of CG, CR, and MINRES in the context of an image deblurring problem [25]. For a given image $\mathbf{x}$, we apply a linear Gaussian smoothing / blurring operator $\mathbf{A}$, as a 2D convolutional operator, on the image and add some noise $\mathbf{e}$ sampled from a uniform distribution, i.e., $\mathbf{b}=\mathbf{A}\mathbf{x}+\mathbf{e}$. We follow a similar experimental setup as the numerical experiment in the work[26]. The true image and blurry version are given in Figure 13. We remind that though $\mathbf{A}$ typically has full rank, it often contains several small singular values, which make it almost numerically singular. We do not use any particular regularization and simply apply CG, CR and MINRES to attempt to recover the original image. The maximum number of iterations and the termination condition are set to $300$ and $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\leq 10^{-5}$, respectively.
The results are presented in Figures 14 and 15. We stress that even when the matrix $\mathbf{A}$ is PD with a few small eigenvalues, it can be nearly numerically singular. Consequently, CG can exhibit behavior similar to that observed in inconsistent systems and may fail to achieve the required termination threshold before reaching the maximum number of iterations. In contrast, both CR and MINRES, as evident from Figures 14 and 15, demonstrate greater stability and terminate once $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\leq 10^{-5}$. Clearly, the image recovered by CG at termination is of significantly lower quality compared to those obtained by CR and MINRES. Similar to the previous section, Figure 15 showcases the best performance of CG — the image recovered by CG at the iteration with the smallest relative residual — compared with the images from CR and MINRES at the same iteration. At this particular iteration, while CG produces a slightly improved image, it remains visibly blurry with a larger relative error than those obtained by CR/MINRES at termination.
@@ 4.2.3 Kernel Regression
Finally, we consider kernel regression [17], which appears frequently in machine learning applications. Kernel regression is an example of non-parametric regression where given training data $\mathcal{D}=\{\mathbf{a}_{i},b_{i}\}_{i=1}^{n}\subset\mathbb{R}^{p}\times\mathbb{R}$, we aim to find a prediction function $f:\mathbb{R}^{p}\to$ that is a solution to the following minimization problem
@@
where $\mathcal{H}$ is some appropriate reproducing kernel Hilbert space (RKHS) with the kernel function $\kappa:^{p}\times^{p}\to$ [32]. It can be shown that any solution to the above problem can be written as $f=g+h$ where $g=\sum_{j=1}^{n}x_{j}\kappa(\mathbf{a}_{j},.)$ and $h\perp\text{Span}\{\kappa(\mathbf{a}_{j},.)\}_{j=1}^{n}$, leading to the least-squares regression problem
@@
We use Wine Quality dataset, comprising $6,497$ training examples with $p=10$ [5]. The dataset is randomly split into $80\%$ training (thus, $n=5,197$) and $20\%$ validation sets. Employing the radial basis function kernel $\kappa(\mathbf{a},\mathbf{a}^{\prime})=\exp{(-L\|\mathbf{a}-\mathbf{a}^{\prime}\|_{2}^{2})}$ with $L=10^{-4}$ results in the positive semi-definite kernel matrix $\mathbf{A}$ of rank $438$. The maximum number of iterations and the termination condition are set to $2,000$ and $\|\mathbf{Ar}_{k}\|/\|\mathbf{Ab}\|\leq 10^{-7}$, respectively. To assess the predictive performance of the recovered model, we measure the mean squared error (MSE) for the test set at each iteration. The results, depicted in Figure 16, clearly show that both CR and MINRES greatly outperform CG, not only in metrics related to the training set but also in terms of predictive performance as measured by the MSE on the test set.
@@ Appendix A Unifying Framework
In this section, we present some technical results for Algorithms 1 and 2 that serve as foundational tools in all of our derivations in this paper. To simplify the exposition, we adopt a more general framework depicted in Algorithm 5, which encapsulates both Algorithms 1 and 2 at the same time. Specifically, considering $s=0$ and $s=1$ in Algorithm 5 amount to Algorithm 1 and Algorithm 2, respectively. In essence, it is shown that all common properties of CG and CR, well-known for positive definite settings, continue to hold under indefinite matrices. The proofs are straightforward adaptation of the PD, found in many classical resources, e.g., [33, 41].
@@
for some $i$. We show the claim for $i+1$. Since $\mathbf{p}_{i-1}\in\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i-1}\mathbf{r}_{0})$, we have $\mathbf{Ap}_{i-1}\in\text{Span}(\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i}\mathbf{r}_{0})$, which from $\mathbf{r}_{i}=\mathbf{r}_{i-1}-\alpha_{i-1}^{(s)}\mathbf{Ap}_{i-1}$ implies $\mathbf{r}_{i}\in\text{Span}(\mathbf{r}_{0},\mathbf{Ar}_{0},\cdots,\mathbf{A}^{i}\mathbf{r}_{0})$. So, $\text{Span}(\mathbf{r}_{0},\mathbf{r}_{1},\cdots,\mathbf{r}_{i})\subset\text{Span}(\mathbf{r}_{0},\cdots,\mathbf{A}^{i}\mathbf{r}_{0})$. However, from $\mathbf{A}^{i}\mathbf{r}_{0}=\mathbf{A}(\mathbf{A}^{i-1}\mathbf{r}_{0})$, we have
@@
by 9a and 10b and our induction assumptions. ∎
Lemma 11.
@@
where the first term and the second term each evaluate to zero by the inductive hypothesis and 9b, respectively. This gives us 11a. Now, let $j=k$, for $i=k$, we have
$\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{p}_{k}\right\rangle=\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\left(\mathbf{r}_{k}+\beta_{k-1}^{(s)}\mathbf{p}_{k-1}\right)\right\rangle=\left\langle\mathbf{r}_{k},\mathbf{A}^{s}\mathbf{r}_{k}\right\rangle,$
by 11a. For $i<k$, we have
@@
where the last line is by the definition of $\beta_{0}^{(s)}$, and the second equality due to 9a and the inductive hypothesis, respectively. This gives us 11b. ∎
@@ Appendix B Further Results
Lemma 12.
Under Assumption 1, in Algorithm 5 we have $(\mathbf{I}-\mathbf{A}^{\dagger}\mathbf{A})\mathbf{p}_{k-1}=c^{(s)}_{k-1}\mathbf{b}^{\perp}$ for $1\leq k\leq g$, where $c^{(s)}_{k-1}=1+\beta^{(s)}_{k-2}c^{(s)}_{k-2},\;k\geq 1$, with $c^{(s)}_{0}=1$.
Proof.